OpenStudy (anonymous):

A purse containing $3.20 in quarters and dimes has, in all, 20 coins. Find the number of each kind of coin.

OpenStudy (cgreenwade2000):

So Quarters (q) + Dimes (d) = 20 and .25q + .1(d) = 3.20

OpenStudy (cgreenwade2000):

From here, we can solve for q or d. q+d=20 So d=20-q

OpenStudy (cgreenwade2000):

Now, we'll go into our second equation and solve it by replacing d with 20-q So our problem looks like .25q +.1(20-q) = 3.20

OpenStudy (cgreenwade2000):

I'll let you solve the rest. But do you see how I set up both equations? The number of quarters is unknown so I called it q. The number of dimes is unknown so I called it d. I know that the number of quarters plus the number of dimes is equal to 20, so my equation became q+d=20 Then, the value of a quarter is .25 so I multiply that by the number of quarters, giving me .25q. Then, the value of a dime is .10 so I multiply that by the number of dimes, giving me .1d. So .25q+.1d=3.20